OpenStudy (anonymous):

Factor each polynomial completely 196w^2-81z^2

OpenStudy (diyadiya):

\[\Large 196w^2-81z^2= (14w)^2-(9z)^2\]

OpenStudy (diyadiya):

\[\Huge a^2-b^2=(a-b)(a+b)\]

OpenStudy (anonymous):

just as ^^ put up difference of squares... a^2 - b^2 = (a+b)(a-b)

OpenStudy (anonymous):

dooh!!

OpenStudy (anonymous):

wow... your LaTeX is even faster than my typing!

OpenStudy (diyadiya):

Haha Thanks :D

OpenStudy (diyadiya):

@Kdilla Did you get the answer ?

OpenStudy (anonymous):

Yes I have a new question now

OpenStudy (anonymous):

What is the binomial factor of 80y^2-120y+45